OpenStudy (anonymous):

If f(x) = −5x + 1, find f(−3). 15 −14 −7 16

Parth (parthkohli):

Plug in -3 wherever you see \(x\).

jhonyy9 (jhonyy9):

f(-3)=-5*(-3) +1 = can you continue ?

jhonyy9 (jhonyy9):

so = 15 +1 =16 right ?